# SteerDiff: Steering towards Safe Text-to-Image Diffusion Models

### v1

@@ 5 Discussion on Concept Forgetting
Figure 5: SteerDiff successfully removes the targeted concepts “Elon Musk” and “Joe Biden” while preserving unrelated concepts such as “Johnny Depp”, “Emma Watson”, “Tom Cruise”, and “Mickey Mouse”. The first 2x2 grid displays the original images generated by SD, while the following three images depict steered samples generated from the same prompt. The image prompts used were: “a photo of X”.
In this section, we explore the potential of SteerDiff to erase specific concepts during image generation. For this experiment, SteerDiff was applied to remove references to “Elon Musk” and “Joe Biden” from user inputs. As shown in Figure 5, the first row illustrates that SteerDiff successfully removed these concepts from the generated images. Notably, SteerDiff effectively removed target concepts while preserving some attributes of the original concepts, such as hairstyle and distinctive clothing style. We also evaluated the method’s effect on unrelated concepts, including “Johnny Depp”, “Emma Watson”, “Tom Cruise”, and “Mickey Mouse”. Ideally, SteerDiff should have little to no impact on unrelated concepts. As outlined in the last two rows, SteerDiff preserved these unrelated concepts. This demonstrates that our approach can be effectively applied to selective concept forgetting without affecting other content.
6 Conclusion
The evolution of diffusion models in generating intricate images highlights both their potential and associated risks. Although recent unlearning methods for diffusion models have made notable progress in mitigating inappropriate content generation, red-teaming studies reveal that these defenses can still be bypassed. Moreover, many defense strategies rely on fine-tuning the model to avoid generating inappropriate content, which becomes increasingly challenging as diffusion models grow larger. In this paper, we present SteerDiff, a lightweight method that acts as an intermediary between the user’s input and the diffusion model, ensuring that generated images comply with ethical and safety standards. We conduct comprehensive experiments across various unlearning concepts to evaluate their effectiveness. Additionally, we benchmark SteerDiff using multiple red-teaming approaches to assess the robustness of our method. Lastly, we explore the potential of SteerDiff in concept removal tasks.
7 Ethics Statement
In this work, we introduce a paradigm to identify inappropriate concepts within input prompts and steer their embeddings to mitigate the generation of unsafe content. Unlike previous methods that primarily focus on post-hoc prevention or concept removal, SteerDiff operates directly on the embeddings of the input prompts, prior to the diffusion process. By intervening earlier in the generation pipeline, we aim to more effectively prevent the propagation of unsafe content.
However, the real-world application of SteerDiff requires carefully defining what constitutes inappropriate concepts, which may vary depending on the application domain. Defining these concepts is a non-trivial task and is likely to require input from human experts, potentially leading to subjective biases. These biases may stem from the social and cultural context in which the system is deployed, as the definition of inappropriateness is highly subjective and dependent on societal norms, which differ across regions and communities. Moreover, since the notion of inappropriateness is largely defined by social norms, the system’s performance may vary depending on whose norms are reflected in the training data. This introduces the risk of reinforcing the biases present in the data, as SteerDiff may disproportionately represent the values of the social groups most prominent in the training set.
Our testbed for evaluating inappropriateness is limited to specific, predefined concepts, which may not fully capture the diversity of opinions and sentiments regarding what is considered inappropriate. As societal norms evolve, so too must the definitions of inappropriateness used by SteerDiff. This necessitates regular updates to the training data and model parameters to ensure the continued relevance and fairness of the system.
Beyond the identification and mitigation of inappropriate content, we believe that SteerDiff can be applied to other areas, such as concept or artistic style removal. As discussed in section 5, SteerDiff has the potential to be extended to various applications, including the removal of specific artistic styles or other undesired concepts in generative models. However, such applications must also carefully consider the ethical implications of content modification, as indiscriminate use of these techniques could lead to censorship or the suppression of artistic expression.
In summary, while SteerDiff offers a promising approach to mitigating unsafe content generation, its reliance on subjective definitions of inappropriateness and the potential for reinforcing societal biases limit its scope and fairness.
8 Reproducibility Statement
Upon acceptance of this paper, all relevant code and data used in our experiments will be made publicly available. The repository will include the source code for SteerDiff, as well as the datasets and instructions necessary to reproduce the results. This will ensure transparency and encourage further research in this domain.
Appendix A Appendix
A.1 Background
A.1.1 Text-to-image diffusion model
Text-to-image (T2I) generation has recently seen significant advancements with the advent of diffusion models (Rombach et al., 2022; 2021; DAL, 2022), which have shown remarkable ability to synthesize high-quality images from textual descriptions. Diffusion models operate by gradually denoising a noisy image, starting from pure Gaussian noise, until a coherent image is formed. The integration of textual information into the diffusion process differentiates T2I diffusion models from traditional diffusion models (Ho & Salimans, 2022), enabling the generation of images that align closely with a given textual input.
In T2I diffusion models, a pre-trained language model, such as CLIP (Radford et al., 2021), is commonly employed to convert text prompts into embeddings. These embeddings are then incorporated into the noise prediction model at various stages of the denoising process. By conditioning the image generation process on these embeddings, the model learns to generate images that not only match the content described in the prompt but also capture finer details of the semantics conveyed by the text. The use of large-scale pre-trained models allows for generalization across a wide range of prompts, enabling the synthesis of highly detailed and contextually appropriate images.
Moreover, recent advances have introduced techniques such as classifier guidance (Ho & Salimans, 2022) and score-based models, which further improve the control over the generated images by adjusting the noise gradients based on the provided text. These methods have made T2I diffusion models not only capable of generating realistic images but also flexible in terms of creative control, making them highly valuable for various generative applications.
A.1.2 Red-teaming text-to-image diffusion models
@@
While T2I diffusion models have demonstrated impressive capabilities, they also pose potential risks related to the generation of harmful or inappropriate content. Red-teaming (Ganguli et al., 2022; Perez et al., 2022; Li et al., 2024), the practice of stress-testing AI systems to uncover vulnerabilities, has become a crucial area of research in ensuring the safety and ethical use of these models. In the context of T2I diffusion models, red-teaming (Chin et al., 2023; Yang et al., 2024; Ma et al., 2024; Zhang et al., 2023b; Zhuang et al., 2023; Tsai et al., 2023) focuses on detecting and mitigating the generation of undesirable or unsafe content, such as offensive imagery or representations of inappropriate concepts.
@@
Given that T2I diffusion models rely heavily on textual prompts to guide image generation, vulnerabilities often arise from the misalignment between the text embeddings and the visual content produced. Adversaries could potentially exploit these models by crafting malicious prompts that subtly bypass content filters. As a result, recent work has explored methods to steer these models away from generating unsafe outputs, such as by applying transformations to the prompt embeddings (Han et al., 2023), or by integrating ethical constraints into the generative process. The challenge lies in maintaining the model’s ability to generate diverse and creative images while ensuring that the output adheres to safety guidelines. Addressing this issue is essential for deploying these models in real-world applications, particularly in sensitive domains such as media, art, and content moderation.
A.2 Training data generation
Following the description in subsubsection 3.1.1, we start from defining a concept set $C_{\text{unsafe}}$ and generate related terms $t_{c}$ entered around one sampled concept $c$ in the blacklist of $C_{\text{unsafe}}$. As introduced in the paper, we do this by prompting an LLM for prompt generation. We follow Latent Guard (Liu et al., 2024; Fan & Tao, 2024; Fan et al., 2024) in conditioning the T2I prompt. For a concept $c$, we prompt an LLM with the following text to generate $t_{c}$:
Prompt for Generating $T_{\text{unsafe}}$: You are a text-to-image diffusion model safety assistant. Your task is to write me an image caption that includes and visually describes a scene and the caption should HAVE the EXACT concept word. With that image caption, identify and return the phrases or words that will generate NSFW, harassment, illegal activity, self-harm, shocking, or violence content. Only return the phrase no other words, no other response.
We also report the LLM prompt used to generate the paired safe phrases $T_{\text{safe}}$ corresponding to above unsafe phrases $T_{\text{unsafe}}$. The prompt is:
Prompt for Generating $T_{\text{safe}}$: You are a NSFW and SFW safety assistant. Return the safe and benign opposite meaning phrase of the given phrase. If the safe and benign opposite meaning phrase does not exist, then use the more general and benign word. Only return the phrase no other words, no other response.
A.3 Red-teaming attack analysis
Figure 6: Attack success rate comparison of SteerDiff and SLD MAX across different categories (lower the better defense performance).
Figure 6 demonstrate the attack success rates (ASR) of two defense mechanisms, SteerDiff and SLD Max, across several categories on the I2P dataset. SteerDiff generally outperforms SLD Max in categories such as harassment, violence, sexual, and shocking, achieving lower ASR values in hate. In contrast, SLD Max demonstrates better defense ability in categories such as hate, self-harm, and illegal activity. Overall, the total ASR suggests a slight advantage for SteerDiff in terms of overall attack vulnerability.
A.4 Visualization of steering
Figure 7 illustrates the outcome of applying the SteerDiff, which learns to project unsafe phrases into a safer latent space. In this 3D PCA visualization, the red points represent the original unsafe phrases, while the green points correspond to the steered versions of these phrases after transformation by SteerDiff.
Figure 7: SteerDiff learns to project unsafe prompts.
A.5 Limitation
In addition to the ethical concerns discussed earlier, SteerDiff has certain technical limitations. One key limitation is the use of a hard-coded steering degree during the experiments, which may lead to over-steering in some cases. This is particularly problematic for prompts containing mildly inappropriate content, where excessive steering could result in unnecessary modifications, thereby altering the intended meaning of the prompt. A more dynamic and context-aware approach could mitigate this issue by adjusting the steering degree based on the severity of the unsafe content, ensuring more precise control without compromising the prompt’s original intent.
A.6 Examples of Generated Images
In our evaluation of different models—SD v1.4, SteerDiff, ESD, SLD STRONG, and SLD MAX. We observe clear trade-offs between the ability to mitigate inappropriate content and the quality of generated images.
SteerDiff consistently demonstrates superior mitigation of inappropriate content while preserving the generative capabilities of the diffusion model. The images produced by SteerDiff maintain higher fidelity, with details and clarity closely resembling those from the baseline SD v1.4, making it the most balanced approach in terms of both safety and image quality. On the other hand, ESD and SLD STRONG offer competitive image quality but occasionally fail to fully suppress inappropriate content generation. These models generate visually appealing images with relatively high quality, especially in complex textures and objects, but their inconsistency in filtering undesirable elements presents a notable shortcoming in safety-critical applications.
SLD MAX achieves a mitigation performance comparable to SteerDiff in terms of defense against inappropriate content generation. However, this comes at the cost of image quality, particularly in depictions of human faces. Images generated with SLD MAX tend to blur facial features, reducing the overall aesthetic quality. This blurring effect is less pronounced in non-human objects but remains a significant limitation in scenarios requiring fine detail preservation.
Overall, SteerDiff stands out as the most effective model for generating high-quality, safe content, whereas SLD MAX provides strong mitigation at the expense of visual detail, particularly in more nuanced areas like facial generation.
A.6.1 Examples of Generated Images in I2P
Figure 8 illustrate the generated example of I2P by Stable Diffusion v1.4, SteerDiff, ESD, SLD STRONG and SLD MAX.
Figure 8: Examples of images generated using I2P Schramowski et al. (2023) prompts. From left to right, the columns represent the augmented prompts, images generated by SteerDiff, ESD, SLD STRONG, and SLD MAX, respectively. Red blocks have been added to the images to obscure explicit inappropriate content.
A.6.2 Examples of Generated Images in COCO-30K
Figure 9 illustrate the generated example of COCO-30k by Stable Diffusion v1.4, SteerDiff, ESD, SLD STRONG and SLD MAX.
Figure 9: Examples of images generated using COCO-30k prompts. From left to right, the columns represent the augmented prompts, images generated by SteerDiff, ESD, SLD STRONG, and SLD MAX, respectively.
A.6.3 Examples of Generated Images Under Sneakyprompt
Figure 10 illustrate the generated example of I2P by SteerDiff, ESD, SLD STRONG and SLD MAX.
Figure 10: Examples of images generated using Sneakyprompt Yang et al. (2024) augmented prompts. From left to right, the columns represent the augmented prompts, and images generated by SteerDiff, ESD, SLD STRONG, and SLD MAX, respectively. Red blocks have been added to the images to obscure explicit inappropriate content.

### v2

@@ 3.6. Artist Style Removal
Figure 3. SteerDiff successfully removes the targeted concepts “Van Gogh” and “Kelly McKernan” while preserving unrelated concepts such as “Pablo Picasso”. The first row displays the original images generated by SD-v1.4, while the second row depicts steered samples generated from the same prompt.
Table 6. Comparison of artist style removal. A higher LPIPSe value indicates more effective target style removal. A lower LPIPSu suggests minimal impact on unrelated artists.
Method Remove “Van Gogh” Remove “Kelly McKernan”
LPIPSe $\uparrow$ LPIPSu $\downarrow$ LPIPSe $\uparrow$ LPIPSu $\downarrow$
CA (Kumari et al., 2023) 0.30 0.13 0.22 0.17
RECE (Gong et al., 2024) 0.31 0.08 0.29 0.04
UCE (Gandikota et al., 2024) 0.25 0.05 0.25 0.03
SteerDiff (Ours) 0.46 0.00 0.47 0.00
We follow the evaluation setup from RECE (Gong et al., 2024; Zhang et al., 2018), where LPIPSe is calculated on the erased artist. A higher LPIPSe value indicates more effective style removal. LPIPSu is calculated on unerased artists, with a lower LPIPSu suggesting minimal impact on unrelated artists. As shown in LABEL:T:dis and qualitative results in Figure 4, SteerDiff achieves superior performance. Our method not only effectively eliminates the target style but also preserves the integrity of non-target styles, by identifying steerable concepts and selectively modifying embeddings only when necessary.
4. Related work
4.1. Attacks on Text-to-Image Models
@@
While T2I diffusion models have demonstrated impressive capabilities, they also pose potential risks related to the generation of harmful or inappropriate content. Red-teaming (Ganguli et al., 2022; Perez et al., 2022; Li et al., 2024; Yang et al., 2024a), the practice of stress-testing AI systems to uncover vulnerabilities, has become a crucial area of research in ensuring the safety and ethical use of these models. In the context of T2I diffusion models, red-teaming (Chin et al., 2023; Yang et al., 2024b; Ma et al., 2024; Zhang et al., 2023; Zhuang et al., 2023; Tsai et al., 2023) focuses on detecting and mitigating the generation of undesirable or unsafe content, such as offensive imagery or representations of inappropriate concepts.
@@
Given that T2I diffusion models rely heavily on textual prompts to guide image generation, vulnerabilities often arise from the misalignment between the text embeddings and the visual content produced. Adversaries could potentially exploit these models by crafting malicious prompts that subtly bypass content filters. As a result, recent works have explored methods to steer these models away from generating unsafe outputs, such as by applying transformations to the prompt embeddings (Han et al., 2023; Li and Beeson, 2025; Yoon et al., 2024; Ahn and Jung, 2025), or by integrating ethical constraints into the generative process. The challenge lies in maintaining the model’s ability to generate diverse and creative images while ensuring that the output adheres to safety guidelines. Addressing this issue is essential for deploying these models in real-world applications, particularly in sensitive domains such as media, art, and content moderation.
4.2. Towards Safe Image Generation
Current approaches to preventing undesirable images from generation generally follow three main strategies. The first involves removing undesired images from the training set, such as excluding all human figures (Nichol et al., 2021) or selectively omitting specific undesirable image categories in the dataset (Schuhmann et al., 2022; DAL, 2022; Rombach and Esser, 2022). However, this approach is costly as it necessitates retraining the model, and removing certain data classes often degrades the overall quality of the generated images (O’Connor, 2022). The second strategy is post-hoc, includes using blacklists for blocking unsafe concepts (DAL, 2022; mid, 2024; Leo, 2023; Markov et al., 2023), model editing methods (Gandikota et al., 2023; Orgad et al., 2023; Gandikota et al., 2024) and unlearning methods(Zhang et al., 2024b; Huang et al., 2024). Blacklists are straightforward to implement, they can be easily bypassed. Model editing methods identify and prune weights within the diffusion model to remove harmful concepts, but they often degrade output quality, thus harming the utility of the models (Rombach et al., 2022). Unlearning methods fine-tune models to erase harmful content, but as model sizes grow, this becomes increasingly computationally expensive and less practical(Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022). Other approaches (Park et al., 2024; Zhang et al., 2024b; Yoon et al., 2024; Gong et al., 2024; Zhang et al., 2024a; Lyu et al., 2024) use inpainting to mask unsafe content or attempt to unlearn inappropriate concepts either within the diffusion model (Zheng and Yeh, 2023) or the textual encoder (Poppi et al., 2023). These methods involve expensive fine-tuning, while SteerDiff offers an intermediary solution that operates between the input prompt and the diffusion model without additional training.
A more promising direction involves training-free, filtering-based techniques. Latent Guard (Liu et al., 2024) identifies inappropriate concepts prior to diffusion by learning a latent space on top of the T2I model’s text encoder; however, it does not actively remove these unsafe representations. In contrast, our proposed method, SteerDiff, acts as an intermediary between the input prompt and the diffusion model, steering unsafe concepts to safe regions without requiring additional model training. Our ablation studies demonstrate that the identifier in SteerDiff substantially outperforms Latent Guard: on the I2P dataset and MS-COCO FID-30K, our identifier achieves classification accuracies of 99.77% and 99.90%, compared to 30.13% and 85.08% for Latent Guard, respectively. Overall, while Latent Guard registers 58.43% accuracy and 30.13% recall, SteerDiff reaches 99.84% accuracy and 99.77% recall (More details are listed in supplementary materials). Similarly, Wu et al. modifies text embeddings by replacing partial embeddings, but our method learns the transformation to steer unsafe concepts to safe while preserving semantics. Moreover, SteerDiff is explicitly designed for safety, incorporating sliding windows for fine-grained unsafe concept detection. This ensures its suitability for safety-driven t2i generation rather than attributes edition.
5. Discussion and Limitation
In this work, we propose SteerDiff, a novel approach that identifies and steers inappropriate concepts in input prompts by intervening directly in the embedding space before the diffusion process begins. However, defining what constitutes “inappropriate” is inherently subjective and varies across cultural and societal contexts, necessitating expert input and raising concerns about bias and fairness. These biases may be amplified by the training data, which often reflects the dominant social norms of specific groups.
In addition to the ethical concerns discussed earlier, SteerDiff has certain technical limitations. SteerDiff currently uses a fixed steering intensity, which may result in overcorrection for borderline cases. A more dynamic and context-aware approach could mitigate this issue by adjusting the steering degree based on the severity of the unsafe content, ensuring more precise control without compromising the original semantics. Our evaluation focuses on a limited set of predefined concepts, suggesting the need for continual updates to remain aligned with evolving norms. Beyond content safety, SteerDiff shows potential for broader applications, such as artistic style removal, though ethical considerations must guide its deployment. In sum, while SteerDiff offers a promising paradigm for safer generative modeling, its effectiveness and fairness hinge on the careful handling of subjectivity, data biases, and ethical considerations.
6. Conclusion
The evolution of diffusion models in generating intricate images highlights both their potential and associated risks. Although recent research for diffusion models has made notable progress in mitigating inappropriate content generation, red-teaming studies reveal that these defenses can still be bypassed. Moreover, many defense strategies rely on fine-tuning the model to avoid generating inappropriate content, which becomes increasingly challenging as diffusion models grow larger. In this paper, we present SteerDiff, a lightweight method that acts as an intermediary between the user’s input and the diffusion model, ensuring that generated images comply with ethical and safety standards. We conduct comprehensive experiments across various unlearning concepts to evaluate their effectiveness. Additionally, we benchmark SteerDiff using multiple red-teaming approaches to assess the robustness of our method. Lastly, we explore the potential of SteerDiff in artist-style removal tasks.